OpenStudy (anonymous):

A cruise ship has circular windows in It's state rooms. What is the circumference of one of these windows If it's diameter is 15 inches ???

OpenStudy (anonymous):

c = Pi * d

OpenStudy (anonymous):

Or it could be written as \(c = \pi d\)

OpenStudy (anonymous):

@gypsy1274 True. Did not want to go through the LaTex process.

OpenStudy (anonymous):

@robtobey Your equation is perfectly clear - I'm just feeling silly.

OpenStudy (anonymous):

@gypsy1274 Don't have any LaTex commands committed to memory. Mathematica is able to copy LaTex and paste the result into a document.

OpenStudy (anonymous):

That would be nice. It is a bit cumbersome to cut and paste equations here.

OpenStudy (anonymous):

@xoxox123 c = circumference and d = diameter. Have you figured out the answer?

OpenStudy (anonymous):

@robtobey Mathematica is expensive, do you use it at home or at work/school?

OpenStudy (anonymous):

Here is an example. Used the mouse to select the expression. Right clicked the mouse and dragged to LaTex. Clicked to place the LaTex code onto the openstudy insert page. clicked the Insert button. \[3 a^2 y \left(a^2 y^2-4 a y+2\right) \]

OpenStudy (anonymous):

Nice. I'll have to stick with \(\LaTeX\) though. My budget doesn't allow for Mathematica.

OpenStudy (anonymous):

On the price of Mathematica. I bought Mathematica 8 Home Edition for $250 as I recall. The agreement is that I cannot use the program for pay. Just upgraded to Mathematica 9 for $95. Can't use it for openstudy because the LaTex feature does't work correctly.

OpenStudy (anonymous):

Does that mean if I purchased the individual license that I could not use it to create worksheets for my students? Or that I can not use it to write books?

OpenStudy (anonymous):

sorry i was eating dinner

OpenStudy (anonymous):

c=3.14(15) ??

OpenStudy (anonymous):

That's ok. Robtobey and I have been having an interesting conversation on your question thread.

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (anonymous):

i got 47.1

OpenStudy (anonymous):

That is correct, unless your teacher wants the answer in terms of \(pi\). Be sure to check your instructions.

OpenStudy (anonymous):

47.1 inches is wrong

OpenStudy (anonymous):

@gypsy1274

OpenStudy (anonymous):

Did you try \(15\pi\)? The calculation is correct and so is the formula.

OpenStudy (anonymous):

Is the diameter 15 or the radius?

OpenStudy (anonymous):

look in the problem it tells you

OpenStudy (anonymous):

and how am i supposed to put that in the answer

OpenStudy (anonymous):

Based on the question you wrote, the answer is 47.1 or \(15\pi\). As for how to input that answer, I have no idea what system or program you are using so I can't answer that. I was asking you to re-check to make sure you had provided accurate information. Sorry, that is the best help I can give you.

OpenStudy (anonymous):

how would i write the second answer

OpenStudy (anonymous):

\(15\pi\)? I don't know how to enter it in your program. Look along the edges of the screen for options find a help screen. You could try the \(\LaTeX\) version, it looks like this: \pi

OpenStudy (anonymous):

soo 15/pi

OpenStudy (anonymous):

i think the website will mark it wrong like that

OpenStudy (anonymous):

No. The slash is going the wrong way.

OpenStudy (anonymous):

This is a question for someone familiar with your program or class. Maybe your teacher?

OpenStudy (anonymous):

okay ..

OpenStudy (anonymous):

could the answer be 15 inches

OpenStudy (anonymous):

Only if your question is asking for the diameter. Based on your question above, 15 is the diameter and not the circumference.

OpenStudy (anonymous):

so whats the circumference ? another way to write 47.1

OpenStudy (anonymous):

Can you provide a screen capture? I might be able to make a better guess if I can see what you are looking at.

OpenStudy (anonymous):

how do i do that ?

OpenStudy (anonymous):

Depends on your computer. A windows computer may have a button labels "scr prt"

OpenStudy (anonymous):

You may also have to use a drawing program like Paint to save the image in .jpg format to upload to this site.